OpenStudy (anonymous):

What is orthonormal basis for R2x2 vector space?

OpenStudy (kinggeorge):

What about\[\left[\begin{matrix} 1&0\\0&0 \end{matrix} \right]\]\[\left[\begin{matrix} 0&1\\0&0 \end{matrix} \right]\]\[\left[\begin{matrix} 0&0\\1&0 \end{matrix} \right]\]\[\left[\begin{matrix} 0&0\\0&1 \end{matrix} \right]\]